OpenStudy (anonymous):

Use substitution to solve this. x+2y=6 3x+6y=8

jimthompson5910 (jim_thompson5910):

first step is to isolate x in `x+2y=6` doing so gives you `x = -2y+6` ------------------------------------- then you move onto the other equation and replace every 'x' you see with '-2y+6' so 3x+6y=8 turns into 3(-2y+6)+6y = 8 notice how the x's are now gone which means you can solve for y. Tell me what you get when you do so

OpenStudy (anonymous):

ok im doing the math now

OpenStudy (anonymous):

y=-4

jimthompson5910 (jim_thompson5910):

what does 3(-2y+6) turn into when you distribute?

OpenStudy (anonymous):

-6y+18

jimthompson5910 (jim_thompson5910):

so 3(-2y+6)+6y = 8 turns into -6y+18+6y = 8

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

on the left side we have -6y and +6y as like terms. They combine to what?

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

so the y terms go away completely leaving us with 18 = 8 but this equation is always false so there are no solutions to this system

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

no problem